OpenStudy (anonymous):

A dehydrated patient needs a 6.4% saline IV. Unfortunately, the hospital only has bags of 6% and 8% saline solutions. How many liters of each of these solutions should be mixed together to yield 1 liter of the desired concentration?

OpenStudy (anonymous):

lets call amount of 8% saline \(x\) then the amount of 6% saline must be \(1-x\) you want \[.08x+.06(1-x)=.064\] or more simply \[8x+(1-x)=64\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

i think i meant \[8x+6(1-x)=6.4\]